# Affine Symmetry and the Group-Theoretic Basis of the Unruh Effect

### v1

@@ IV Inequivalent Minkowski and Rindler modes in one-dimensional affine group representations
$b_{\pm}(\omega,\lambda)=\frac{1}{\sqrt{2\pi}}\int^{+\infty}_{0}\frac{dk}{k}k^{\mp i(\omega-i\lambda)}a(k)\,.$ (62)
@@
$c_{+}(\omega)=\frac{1}{\sqrt{2\pi}}\Gamma(-i\omega)[e^{-\frac{\pi}{2}\omega}\ b_{+}^{\dagger}(\omega)+e^{+\frac{\pi}{2}\omega}\ b_{-}(\omega)]$ (66)
@@
$|\omega\rangle=\frac{1}{\sqrt{2\pi}}\int_{0}^{\infty}\frac{dk}{k}\,k^{i\omega}\,|k\rangle_{+},\qquad\langle\omega|\omega^{\prime}\rangle=\delta(\omega-\omega^{\prime})\,.$ (68)
@@
$_{+}\langle\xi|\omega\rangle=\frac{1}{\sqrt{2\pi}}\int_{0}^{\infty}\frac{dk}{k}\,k^{i\omega}e^{ike^{\xi}}=\frac{1}{\sqrt{2\pi}}\,e^{i\omega\xi}\,e^{-\frac{\pi\omega}{2}}\Gamma(i\omega),$ (73)
@@
$\langle\omega|\omega^{\prime}\rangle=|\Gamma(i\omega)|^{2}\,e^{-\pi\omega}\,\delta(\omega-\omega^{\prime})=\frac{2\pi}{\omega}\,\frac{1}{e^{2\pi\omega}-1}\,\delta(\omega-\omega^{\prime}),$ (74)
where we used $|\Gamma(i\omega)|^{2}=\pi/[\omega\sinh(\pi\omega)]$. This weighted measure differs from the flat measure defining the natural inner product (71) on $L^{2}(\mathbb{R},d\xi)$, with respect to which the plane waves $|\omega\rangle_{\xi}$ are orthonormal. Consequently, the map

### v2

@@ IV Inequivalent Minkowski and Rindler modes in one-dimensional affine group representations
$b_{\pm}(\omega,\lambda)=\frac{1}{\sqrt{2\pi}}\int^{+\infty}_{0}\frac{dk}{k}k^{{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\pm}i(\omega-i\lambda)}a(k)\,.$ (62)
@@
$c_{+}(\omega)=\frac{1}{\sqrt{2\pi}}\Gamma({\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}+}i\omega)[e^{-\frac{\pi}{2}\omega}\ b_{{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}-}}^{\dagger}(\omega)+e^{+\frac{\pi}{2}\omega}\ b_{{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}+}}(\omega)]$ (66)
@@
$|\omega\rangle=\frac{1}{\sqrt{2\pi}}\int_{0}^{\infty}\frac{dk}{k}\,k^{{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}-}i\omega}\,|k\rangle_{+},\qquad\langle\omega|\omega^{\prime}\rangle=\delta(\omega-\omega^{\prime})\,.$ (68)
@@
$_{+}\langle\xi|\omega\rangle=\frac{1}{\sqrt{2\pi}}\int_{0}^{\infty}\frac{dk}{k}\,k^{{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}-}i\omega}e^{ike^{\xi}}=\frac{1}{\sqrt{2\pi}}\,e^{i\omega\xi}\,e^{{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}+}\frac{\pi\omega}{2}}\Gamma({\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}-}i\omega),$ (73)
@@
$\langle\omega|\omega^{\prime}\rangle=|\Gamma({\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}-}i\omega)|^{2}\,e^{\pi\omega}\,\delta(\omega-\omega^{\prime}).$ (74)
The additional $\omega$-dependent factor on the right hand side renders this expression different from the plain $\delta$-function defining the natural inner product (71) on $L^{2}(\mathbb{R},d\xi)$, with respect to which the plane waves $|\omega\rangle_{\xi}$ are orthonormal. Consequently, the map